Feb 24, 2007. 

 

The last 24 hours have seen an amazing amount of accomplishments on the five probes. All five EFI and SCM instruments were turned on, a calibration of Probe B FGM was performed, and all probes were spun down to 11 RPM.  We recovered all engineering and science data from all probes, too.

 

As the probes have begun to separate a little, we were also able to contact three probes simultaneously using three different ground stations.  Probes E, D and C telemetered to Berkeley, Santiago and Wallops at 512, 256 and 128 kbps, and the data routing network worked perfectly.  Of the 20 passes during the day, Berkeley had 13, Santiago 1, Haartebeestok 1 and Wallops had 5 tracking passes.

 

All five probes remain in excellent health.  Temperatures are very mild on the probes now in their nominal attitude with sun on their side panels.